OpenStudy (anonymous):

2nd. an athletics track is to have a perimeter of 400m and straights of 78m. Find the radius of the circular bond

OpenStudy (mertsj):

What do you mean by "circular bond"?

OpenStudy (anonymous):

\[(400-2*78)/(2\pi)=122/\pi=38.8338...\]